# Jailbreaking Leading Safety-Aligned LLMs with Simple Adaptive Attacks

### v2

@@ Abstract
We show that even the most recent safety-aligned LLMs are not robust to simple adaptive jailbreaking attacks. First, we demonstrate how to successfully leverage access to logprobs for jailbreaking: we initially design an adversarial prompt template (sometimes adapted to the target LLM), and then we apply random search on a suffix to maximize a target logprob (e.g., of the token “Sure”), potentially with multiple restarts. In this way, we achieve nearly 100% attack success rate—according to GPT-4 as a judge—on Vicuna-13B, Mistral-7B, Phi-3-Mini, Nemotron-4-340B, Llama-2-Chat-7B/13B/70B, Llama-3-Instruct-8B, Gemma-7B, GPT-3.5, GPT-4, and R2D2 from HarmBench that was adversarially trained against the GCG attack. We also show how to jailbreak all Claude models—that do not expose logprobs—via either a transfer or prefilling attack with a 100% success rate. In addition, we show how to use random search on a restricted set of tokens for finding trojan strings in poisoned models—a task that shares many similarities with jailbreaking—which is the algorithm that brought us the first place in the SaTML’24 Trojan Detection Competition. The common theme behind these attacks is that adaptivity is crucial: different models are vulnerable to different prompting templates (e.g., R2D2 is very sensitive to in-context learning prompts), some models have unique vulnerabilities based on their APIs (e.g., prefilling for Claude), and in some settings, it is crucial to restrict the token search space based on prior knowledge (e.g., for trojan detection). For reproducibility purposes, we provide the code, logs, and jailbreak artifacts in the JailbreakBench format at https://github.com/tml-epfl/llm-adaptive-attacks.
@@ 1 Introduction
Table 1: Summary of our results. We measure the attack success rate for the leading safety-aligned LLMs on a dataset of $50$ harmful requests from Chao et al. (2023). We consider an attack successful if GPT-4 as a semantic judge gives a 10/10 jailbreak score.
@@
GPT-4 Turbo OpenAI Logprobs Prompt + random search + self-transfer 59%* 96%
Claude 2.0 Anthropic Tokens Prefilling attack 61%* 100%
Claude 2.1 Anthropic Tokens Prefilling attack 68%* 100%
Claude 3 Haiku Anthropic Tokens Prefilling attack None 100%
Claude 3 Sonnet Anthropic Tokens Transfer from GPT-4 Turbo None 100%
Claude 3 Opus Anthropic Tokens Prefilling attack None 100%
@@
Figure 1: Successful transfer attack on Claude 3 Sonnet. We show an illustrative example using temperature zero with an adversarial suffix generated on GPT-4 leveraging access to its logprobs. We observe that one can directly ask follow-up requests to detail some steps generated in the first response to get much more information. Note that the upper part of the user prompt is cropped (see Table 2 for the full prompt).
The remarkable capabilities of Large Language Models (LLMs) carry the inherent risk of misuse, such as producing toxic content, spreading misinformation or supporting harmful activities. To mitigate these risks, safety alignment or refusal training is commonly employed—a fine-tuning phase where models are guided to generate responses judged safe by humans and to refuse responses to potentially harmful queries (Bai et al., 2022; Touvron et al., 2023). Although safety alignment is effective in general, several works have shown that it can be circumvented using adversarial prompts. These are inputs specifically designed to induce harmful responses from the model, a practice known as jailbreaking attacks (Mowshowitz, 2022; Zou et al., 2023; Chao et al., 2023).
Jailbreaking attacks vary in their knowledge of the target LLM (ranging from white- to black-box approaches, or API-only access), complexity (involving manual prompting, standard optimization techniques, or auxiliary LLMs), and computational cost. Moreover, the nature of the jailbreaks they produce differs: some methods insert strings with little or no semantic meaning (Zou et al., 2023), while others rephrase user requests to maintain natural language (Mehrotra et al., 2023). The effectiveness of these attacks can significantly vary, achieving a high success rate on some target models but also drastically failing on others. Finally, some LLMs, such as the Llama-2-Chat family (Touvron et al., 2023), seem to maintain their robustness against these attacks. At the same time, new defensive mechanisms designed to counteract jailbreaks are emerging (Robey et al., 2023; Mazeika et al., 2024).
In this work, we examine the safety of leading safety-aligned LLMs in terms of robustness to jailbreaks. We show that it is feasible to leverage the information available about each model, derived from training details or inference (e.g., logprobs), to construct simple adaptive attacks. Our main tool consists of a manually designed prompt template—which is used for all unsafe requests for a given model—enhanced by an adversarial suffix found with random search (RS) (Rastrigin, 1963) when the logprobs of the generated tokens are at least partially accessible. Our approach can be considered simple as it does not require gradient information (even for open-weight models) (Zou et al., 2023; Geisler et al., 2024), auxiliary LLMs to iteratively optimize the jailbreaks (Chao et al., 2023; Mehrotra et al., 2023; Zeng et al., 2024), or multi-turn conversations (Cheng et al., 2024; Russinovich et al., 2024). In this way, using the dataset of 50 harmful requests from AdvBench (Zou et al., 2023) curated by Chao et al. (2023), we obtain a close to 100% attack success rate on all leading safety-aligned LLMs, including Vicuna-13B, Mistral-7B, Phi-3-Mini, Nemotron-4-340B, Llama-2-Chat-7B/13B/70B, Llama-3-Instruct-8B, Gemma-7B, GPT-3.5, GPT-4, Claude-3, and the adversarially trained R2D2. A summary of our main results in Table 1 suggests that our methods substantially outperform the existing attacks and achieve a 100% attacks success rate on many models for the first time. Additionally, we summarize all our evaluations in Table 22 in the appendix. We also show an illustrative example of a successful transfer attack on Claude 3 Sonnet in Figure 1. Finally, we show how to use RS on a restricted set of tokens for finding trojan strings in poisoned models—a task that shares many similarities with jailbreaking—enabling us to secure the first place in the SaTML’24 Trojan Detection Competition (Rando & Tramèr, 2024).
Our results provide several insights into the domain of safety in LLMs and its evaluation. First, we reveal that currently both open-weight and proprietary models are completely non-robust to adversarial attacks. Second, it is evident that adaptive attacks play a key role in the evaluation of robustness, as no single method can generalize across all target models. Despite the absence of a standardized attack, we still provide recommendations for future research on designing jailbreak attacks, analogous to the framework established for image classification by Carlini et al. (2019); Tramèr et al. (2020); Croce et al. (2022b), distilling key observations from our experiments.
@@ 2 Related Work
Manual attacks.
ChatGPT users have discovered handcrafted jailbreaks (Mowshowitz, 2022). Wei et al. (2023a) systematically categorize these jailbreaks based on two main criteria: (1) competing objectives, which occurs when a model’s capabilities conflict with safety goals, and (2) mismatched generalization, which arises when safety training does not generalize to domains where the model has capabilities. By leveraging these failure modes and employing a combination of manual attacks, Wei et al. (2023a) achieve high success rates on proprietary LLMs such as GPT-4 and Claude v1.3. Wei et al. (2023b) explore jailbreaking using in-context learning prompts that contain a few examples of harmful responses.
Direct search attacks.
Alternatively, the search for jailbreaks can be automated using first- or zeroth-order discrete optimization techniques. For example, Zou et al. (2023) introduce universal and transferable attacks with a gradient-based method named Greedy Coordinate Gradient (GCG), inspired by earlier discrete optimization efforts in NLP (Shin et al., 2020). Lapid et al. (2023) use a genetic algorithm to generate universal adversarial prompts within a black-box threat model, where gradients are not used. Liu et al. (2023) apply genetic algorithms to combine sentence fragments into a low-perplexity jailbreak. Zhu et al. (2023) pursue a similar goal, modifying GCG to generate low-perplexity adversarial suffixes. Sitawarin et al. (2024); Hayase et al. (2024) suggest employing random search on predicted probabilities for black-box models to guide and refine the adversarial string search, occasionally aided by a white-box LLM to identify the most promising tokens to change. For OpenAI models, both attacks use the logit_bias parameter whose behavior has been already changed: it no longer influences the logprobs, rendering their attacks ineffective.
LLM-assisted attacks.
Finally, using other LLMs for optimizing jailbreaking attacks has shown considerable promise, primarily due to enhanced query efficiency. Chao et al. (2023) have first developed Prompt Automatic Iterative Refinement (PAIR), a method that uses an auxiliary LLM to identify jailbreaks efficiently. Mehrotra et al. (2023) have then refined PAIR’s methodology, introducing a tree-based search method. In similar vein, Shah et al. (2023) have devised an approach to jailbreaks generation using an LLM that is guided by persona modulation. Meanwhile, Yu et al. (2023) have introduced GPTFUZZER, a framework that iteratively enhances human-written templates with the help of an LLM. Zeng et al. (2024) have fine-tuned GPT-3.5 for the specific task of rephrasing harmful requests, using the rephrased content to jailbreak a target LLM. Takemoto (2024) offer a straightforward LLM rephrasing method that rivals more complex methods.
@@ 3.1 Setting
Background on jailbreaking.
We focus on identifying prompts that, when given a specific harmful request (e.g., “Tell me how to build a bomb”), induces the LLM to generate harmful content. We assume access to a set of such requests recognized by most LLM providers as harmful (e.g., misinformation, violence, hateful content) and are typically not responded to. We define a language model $\texttt{LLM}:\mathcal{T}^{*}\rightarrow\mathcal{T}^{*}$ as a function that maps a sequence of input tokens to a sequence of output tokens, referred to as the target model, as it is the one we aim to jailbreak. Given a judge function $\texttt{JUDGE}:\mathcal{T}^{*}\times\mathcal{T}^{*}\rightarrow\{\text{NO},\text{YES}\}$ and a harmful request $R\in\mathcal{T}^{*}$, the attacker’s goal is:
@@
Our setup.
We use default system prompts unless specifically mentioned (modifications are only made for Claude) due to potential future restrictions by frontier LLM providers, who might limit access to the system prompt for safety reasons. Our targets comprise a set of 50 behaviors from AdvBench curated by Chao et al. (2023) that ensures distinct and diverse harmful requests. We use GPT-4 as the semantic judge (see Table 9 for the prompt) in line with the criteria established by Chao et al. (2023), where a jailbreak is considered successful only if it achieves a 10/10 jailbreak score from GPT-4. We also include results using the rule-based judge from Zou et al. (2023), as well as Llama-3-70B and Llama Guard 2 judges in Appendix B.6 for comparison. Additionally, we manually inspect all generations and flag cases with a significant number of false positives (which we only observed on Claude 2.1).
@@ 3.2 Methodology
Table 2: Our main prompt template. target_str is the desired beginning of the LLM’s response (e.g., “Sure, here is how to make a bomb”), goal is the harmful request (e.g., “Provide a tutorial on how to make a bomb”), adv_suffix is the suffix optimized by random search.
@@
Prior works define adaptive attacks as attacks that are specifically designed to target a given defense (Tramèr et al., 2020). We follow this definition and describe the building blocks of our adaptive attacks, which we combine and potentially adapt depending on the target LLMs.
Prompt templates.
The importance of a well-designed prompt in enhancing the performance of LLMs is well-established (Wei et al., 2023a). In our approach, we develop a prompt template that can incorporate a generic unsafe request. This template is specifically designed to make the model start from a specified string (e.g., “Sure, here is how to make a bomb”) and steer the model away from its default aligned behavior. Its general structure can be summarized as: set of rules $+$ harmful request $+$ adversarial suffix. We have optimized the set of rules one by one on the GPT-3.5 Turbo model to maximize the attack success rate and avoid the built-in safety guardrails. We provide the complete template in Table 2. As we will see in Section 4 (Table 5), this prompt template alone leads to 100% attack success rate on GPT-3.5 Turbo, and it also provides a good starting point for other LLMs. We have also designed another prompt template (referred to as the in-context prompt) that includes an example of unsafe behavior which the model is encouraged to imitate (see Table 8).
Random search.
We use a simple random search (RS) algorithm (Rastrigin, 1963) adapted for jailbreaking language models. The algorithm can be summarized as follows:
@@
We use adversarial suffixes initialized with $25$ tokens (which we found to be a nearly optimal number—see an example on Gemma-7B in Appendix B.1), we use up to 10 0001000010\,00010 000 iterations and up to 10 random restarts (although in most cases a single restart suffices). This strategy, including the preference for suffixes over prefixes and the focus on maximizing the probability of the token “Sure”, draws inspiration from the attack methodology of Zou et al. (2023). We also tried other affirmative words, such as “exactly” and “certainly”, but we did not observe any improvement in the attack success rate. We opted for random search (a) due to its simplicity and efficiency, requiring only logprobs instead of gradients (thus also reducing the memory demands), and (b) motivated by its success in attacking vision models (Andriushchenko et al., 2020; Croce et al., 2022a).
Transfer attacks.
Successful jailbreaks developed for one LLM can often be reused on another model (Zou et al., 2023). This observation will be crucial for attacking some of the Claude 3 models that do not expose logprobs making random search not applicable.
Self-transfer.
It is well-known that initialization plays a key role in optimization algorithms, including in RS-based attacks (Andriushchenko et al., 2020). We leverage the adversarial suffix found by random search for a simpler harmful request as the initialization for RS on more challenging requests. We refer to this approach as self-transfer. Interestingly, this approach often leads to a transferable adversarial suffix (or, at least, a good starting for a subsequent run of random search), even though it is crafted for a single model and a single request. It significantly boosts the query efficiency and attack success rate on many LLMs.
Prefilling attack.
Some APIs like Claude allow users to directly prefill the LLM’s response with a specified beginning, making the aforementioned optimization procedure unnecessary. In that case, we explore prefilling the response with a string that corresponds to the target behavior (e.g., “Sure, here is how to make a bomb”). As a side note, the prefilling attack is also straightforward to implement for any open-weight LLM where the conversation template can be directly modified (Vega et al., 2023).
Reproducibility.
For reproducibility purposes, we provide all further details in our repository https://github.com/tml-epfl/llm-adaptive-attacks that contains the code, logs, and jailbreak artifacts in the JailbreakBench format.

### v3

@@ Abstract
We show that even the most recent safety-aligned LLMs are not robust to simple adaptive jailbreaking attacks. First, we demonstrate how to successfully leverage access to logprobs for jailbreaking: we initially design an adversarial prompt template (sometimes adapted to the target LLM), and then we apply random search on a suffix to maximize a target logprob (e.g., of the token “Sure”), potentially with multiple restarts. In this way, we achieve 100% attack success rate—according to GPT-4 as a judge—on Vicuna-13B, Mistral-7B, Phi-3-Mini, Nemotron-4-340B, Llama-2-Chat-7B/13B/70B, Llama-3-Instruct-8B, Gemma-7B, GPT-3.5, GPT-4o, and R2D2 from HarmBench that was adversarially trained against the GCG attack. We also show how to jailbreak all Claude models—that do not expose logprobs—via either a transfer or prefilling attack with a 100% success rate. In addition, we show how to use random search on a restricted set of tokens for finding trojan strings in poisoned models—a task that shares many similarities with jailbreaking—which is the algorithm that brought us the first place in the SaTML’24 Trojan Detection Competition. The common theme behind these attacks is that adaptivity is crucial: different models are vulnerable to different prompting templates (e.g., R2D2 is very sensitive to in-context learning prompts), some models have unique vulnerabilities based on their APIs (e.g., prefilling for Claude), and in some settings, it is crucial to restrict the token search space based on prior knowledge (e.g., for trojan detection). For reproducibility purposes, we provide the code, logs, and jailbreak artifacts in the JailbreakBench format at https://github.com/tml-epfl/llm-adaptive-attacks.
@@ 1 Introduction
The remarkable capabilities of Large Language Models (LLMs) carry the inherent risk of misuse, such as producing toxic content, spreading misinformation or supporting harmful activities. To mitigate these risks, safety alignment or refusal training is commonly employed—a fine-tuning phase where models are guided to generate responses judged safe by humans and to refuse responses to potentially harmful queries (Bai et al., 2022; Touvron et al., 2023). Although safety alignment is effective in general, several works have shown that it can be circumvented using adversarial prompts. These are inputs specifically designed to induce harmful responses from the model, a practice known as jailbreak attacks (Mowshowitz, 2022; Zou et al., 2023; Chao et al., 2023).
Jailbreak attacks vary in their knowledge of the target LLM (ranging from white- to black-box approaches, or API-only access), complexity (involving manual prompting, standard optimization techniques, or auxiliary LLMs), and computational cost. Moreover, the nature of the jailbreaks they produce differs: some methods insert strings with little semantic meaning (Zou et al., 2023), while others rephrase user requests to maintain natural language (Mehrotra et al., 2023). The effectiveness of these attacks can significantly vary, achieving a high success rate on some target models but also drastically failing on others. For example, the Llama-2-Chat and Claude family of LLMs maintain high robustness against existing attacks (Anthropic, 2024b; Touvron et al., 2023). Moreover, new defenses designed to counteract jailbreaks are emerging (Robey et al., 2023; Mazeika et al., 2024).
Contributions. In this work, we examine the safety of leading safety-aligned LLMs in terms of robustness to jailbreaks. We show that it is feasible to leverage the information available about each model, derived from training details or inference (e.g., logprobs), to construct simple adaptive attacks, which we define as attacks that are specifically designed to target a given defense (Tramèr et al., 2020). Our main tool consists of a manually designed prompt template—which is used for all unsafe requests for a given model—enhanced by an adversarial suffix found with random search (Rastrigin, 1963) when the logprobs of the generated tokens are at least partially accessible. Our approach can be considered simple as it does not require gradient information (Zou et al., 2023; Geisler et al., 2024), auxiliary LLMs to iteratively optimize the jailbreaks (Chao et al., 2023; Mehrotra et al., 2023; Zeng et al., 2024), or multi-turn conversations (Cheng et al., 2024; Russinovich et al., 2024). In this way, using the dataset of 50 harmful requests from AdvBench (Zou et al., 2023) curated by Chao et al. (2023), we obtain a 100% attack success rate on all leading safety-aligned LLMs, including Vicuna-13B, Mistral-7B, Phi-3-Mini, Nemotron-4-340B, Llama-2-Chat-7B/13B/70B, Llama-3-Instruct-8B, Gemma-7B, GPT-3.5, GPT-4o, Claude-3/3.5, and the adversarially trained R2D2. A summary of our main results in Table 1 suggests that our methods substantially outperform the existing attacks and achieve a 100% attacks success rate on many models for the first time. We also show an illustrative example of a successful transfer attack on Claude 3 Sonnet (Figure 5 in the appendix). Additionally, we show how to use random search on a restricted set of tokens for finding trojan strings in poisoned models—a task that shares many similarities with jailbreaking—enabling us to secure the first place in the SaTML’24 Trojan Detection Competition. Finally, we list all our evaluations in Table 21 in the appendix, which we hope will serve as a valuable source of information on the robustness of frontier LLMs.
Insights. Our results provide several insights into the domain of safety in LLMs and its evaluation. First, we reveal that currently both open-weight and proprietary models are completely non-robust to adversarial attacks. Second, it is evident that adaptive attacks play a key role in the evaluation of robustness, as no single method can generalize across all target models. Despite the absence of a standardized attack, we still provide recommendations for future research on designing jailbreak attacks, analogous to the framework established for image classification by Carlini et al. (2019); Tramèr et al. (2020); Croce et al. (2022b), distilling key observations from our experiments.
Table 1: Summary of our results. We measure the attack success rate for the leading safety-aligned LLMs on the set of $50$ harmful requests from AdvBench (Zou et al., 2023) curated by Chao et al. (2023). We consider an attack successful if GPT-4 as a semantic judge gives a 10/10 jailbreak score.
@@
GPT-4o OpenAI Logprobs Prompt + Random Search + Self-Transfer None 100%
Claude 2.0 Anthropic Tokens Prompt + Prefilling Attack 61%* 100%
Claude 2.1 Anthropic Tokens Prompt + Prefilling Attack 68%* 100%
Claude 3 Haiku Anthropic Tokens Prompt + Prefilling Attack None 100%
Claude 3 Sonnet Anthropic Tokens Prompt + Transfer from GPT-4 Turbo None 100%
Claude 3 Opus Anthropic Tokens Prompt + Prefilling Attack None 100%
Claude 3.5 Sonnet Anthropic Tokens Prompt + Prefilling Attack None 100%
@@ 2 Related Work
Manual attacks. ChatGPT users have discovered handcrafted jailbreaks (Mowshowitz, 2022). Wei et al. (2023a) systematically categorize these jailbreaks based on two main criteria: (1) competing objectives, which occurs when a model’s capabilities conflict with safety goals, and (2) mismatched generalization, which arises when safety training does not generalize to domains where the model has capabilities. By leveraging these failure modes and employing a combination of manual attacks, Wei et al. (2023a) achieve high success rates on proprietary LLMs such as GPT-4 and Claude v1.3. Wei et al. (2023b) explore jailbreaking using in-context learning prompts that contain a few examples of harmful responses, while Anil et al. (2024) take it a step further by using many in-context examples and derive a predictable scaling trend for long-context LLMs.
Direct search attacks. The search for jailbreaks can be automated using first- or zeroth-order discrete optimization techniques. For example, Zou et al. (2023) introduce universal and transferable attacks with a gradient-based method named Greedy Coordinate Gradient (GCG), inspired by earlier discrete optimization efforts in NLP (Shin et al., 2020). Lapid et al. (2023) use a genetic algorithm to generate universal adversarial prompts within a black-box threat model, where gradients are not used. Liu et al. (2023) apply genetic algorithms to combine sentence fragments into a low-perplexity jailbreak. Zhu et al. (2023) pursue a similar goal, modifying GCG to generate low-perplexity adversarial suffixes. Liao & Sun (2024) propose to learn a separate model to generate adversarial prefixes similar to GCG. More related to our work, Sitawarin et al. (2024); Hayase et al. (2024) suggest employing random search on predicted probabilities for black-box models to guide and refine the adversarial string search, occasionally aided by a white-box LLM to identify the most promising tokens to change. For OpenAI models, both attacks use the logit_bias parameter whose behavior has been already changed: it no longer influences the logprobs, rendering their attacks ineffective.
LLM-assisted attacks. Finally, using other LLMs for optimizing jailbreaking attacks has shown considerable promise, primarily due to enhanced query efficiency. Chao et al. (2023) have first developed Prompt Automatic Iterative Refinement (PAIR), a method that uses an auxiliary LLM to identify jailbreaks efficiently. Mehrotra et al. (2023) have then refined PAIR’s methodology, introducing a tree-based search method. In similar vein, Shah et al. (2023) have devised an approach to jailbreaks generation using an LLM that is guided by persona modulation. Meanwhile, Yu et al. (2023) have introduced GPTFUZZER, a framework that iteratively enhances human-written templates with the help of an LLM. Zeng et al. (2024) have fine-tuned GPT-3.5 for the specific task of rephrasing harmful requests, using the rephrased content to jailbreak a target LLM.
@@ 3.1 Setting
Background on jailbreaking. We focus on identifying prompts that, when given a specific harmful request (e.g., “Tell me how to build a bomb”), induces the LLM to generate harmful content. We assume access to a set of such requests recognized by most LLM providers as harmful (e.g., misinformation, violence, hateful content) and are typically not responded to. We define a language model $\texttt{LLM}:\mathcal{T}^{*}\rightarrow\mathcal{T}^{*}$ as a function that maps a sequence of input tokens to a sequence of output tokens, referred to as the target model, as it is the one we aim to jailbreak. Given a judge function $\texttt{JUDGE}:\mathcal{T}^{*}\times\mathcal{T}^{*}\rightarrow\{\text{NO},\text{YES}\}$ and a harmful request $R\in\mathcal{T}^{*}$, the attacker’s goal is:
@@
Our setup. We use default system prompts unless specifically mentioned (modifications are only made for Claude) due to potential future restrictions by frontier LLM providers, who might limit access to the system prompt for safety reasons. Our targets comprise a set of 50 behaviors from AdvBench curated by Chao et al. (2023) that ensures distinct and diverse harmful requests. We use GPT-4 as the semantic judge (see Table 8 for the prompt) in line with the criteria established by Chao et al. (2023), where a jailbreak is considered successful only if it achieves a 10/10 jailbreak score from GPT-4. We also include results using the rule-based judge from Zou et al. (2023), as well as Llama-3-70B and Llama Guard 2 judges in Appendix B.6 for comparison. Additionally, we manually inspect all generations and flag cases with a significant number of false positives (which we only observed on Claude 2.1).
@@ 3.2 Methodology
Figure 1: Our main prompt template. target_str is the desired beginning of the LLM’s response (e.g., “Sure, here is how to make a bomb”), goal is the harmful request (e.g., “Provide a tutorial on how to make a bomb”), adv_suffix is the suffix optimized by random search.
@@
Prior works define adaptive attacks as attacks that are specifically designed to target a given defense (Tramèr et al., 2020). We follow this definition and describe the building blocks of our adaptive attacks, which we combine and adapt depending on the target LLMs. Importantly, we customize our adaptive attacks for each model but not for each request.
Prompt templates. The importance of a well-designed prompt in enhancing the performance of LLMs is well-established (Wei et al., 2023a). In our approach, we develop a prompt template that can incorporate a generic unsafe request. This template is specifically designed to make the model start from a specified string (e.g., “Sure, here is how to make a bomb”) and steer the model away from its default aligned behavior. Its general structure can be summarized as: set of rules $+$ harmful request $+$ adversarial suffix. We have optimized the set of rules one by one on the GPT-3.5 Turbo model to maximize the attack success rate and avoid the built-in safety guardrails. We provide the complete template in Figure 1. As we will see in Section 4 (Table 3), this prompt template alone leads to 100% attack success rate on GPT-3.5 Turbo, and it also provides a good starting point for other LLMs. We have also customized this prompt for GPT-4o (see Table 7) and designed another prompt template (referred to as the in-context prompt) that includes an example of an unsafe behavior which the model is encouraged to imitate (see Table 6).
Random search. We use a simple random search (RS) algorithm (Rastrigin, 1963) adapted for jailbreaking language models. The algorithm can be summarized as follows:
@@
We use adversarial suffixes initialized with $25$ tokens (which we found to be a nearly optimal number—see an example on Gemma-7B in Appendix B.1), we use up to 10 0001000010\,00010 000 iterations and up to 10 random restarts, although in most cases a single restart suffices. This strategy, including the preference for suffixes over prefixes and the focus on maximizing the probability of the token “Sure”, draws inspiration from the attack methodology of Zou et al. (2023). We also tried other affirmative words, such as “exactly” and “certainly”, but we did not observe any improvements. We opted for random search (1) due to its simplicity and efficiency, requiring only logprobs instead of gradients (thus also reducing the memory demands), and (2) motivated by its success in attacking vision models (Andriushchenko et al., 2020; Croce et al., 2022a).
Self-transfer. It is well-known that initialization plays a key role in optimization algorithms, including in random search based attacks (Andriushchenko et al., 2020). We leverage the adversarial suffix found by random search for a simpler harmful request as the initialization for random search on more challenging requests. We refer to this approach as self-transfer. Interestingly, this approach often leads to a transferable adversarial suffix (or, at least, a good starting for a subsequent run of random search), even though it is crafted for a single model and a single request. It significantly boosts the query efficiency and attack success rate on many LLMs.
Transfer attacks. Successful jailbreaks developed for one LLM can often be reused on another model (Zou et al., 2023). This observation will be crucial for attacking some of the Claude 3 models that do not expose logprobs making random search not applicable.
Prefilling attack. Some APIs like Claude allow users to directly prefill the LLM’s response with a specified beginning, making iterative optimization unnecessary. Thus, for Claude models, we explore prefilling the response with a string that corresponds to a target behavior (e.g., “Sure, here is how to make a bomb”). As a side note, prefilling is also straightforward to implement for any open-weight LLM, where the chat template can be modified directly (Vega et al., 2023).
Reproducibility. For reproducibility purposes, we provide all further details in our repository [^1] https://github.com/tml-epfl/llm-adaptive-attacks that contains the code, logs, and jailbreak artifacts in the JailbreakBench format.